Two uniform metal spheres with masses of 100kg and 150kg have radii of 6.0 cm and 8.0 cm, respectively. If their surfaces are separated by a distance of 0.86 m, what is the gravitational force of attraction between the spheres?
Answer #1

Gravitational force between two objects is given by:

F= G*m1*m2/R^2

m1 = 100 kg, m2 = 150 kg

R = distance between centers of both object = r1 + r2 + d

R = 0.06 + 0.08 + 0.86 = 1.00 m

So,

F = 6.67*10^-11*100*150/1.00^2

F = 1.00*10^-6 N

Please Upvote.